Prove that a non-decreasing (resp. non-increasing) sequence which is not bounded above (resp. bounded below) diverges to +∞ (resp. to −∞).


1
Expert's answer
2021-02-03T15:17:31-0500

1) Let "\{a_n\}_{n=1}^{+\infty}"  is a non-decreasing sequence which is not bounded above, that is, "\sup\limits_{n}a_n = +\infty". This means that for all M>0 there exists "N\in\mathbb{N}" such that "a_N>M". As the sequence "\{a_n\}_{n=1}^{+\infty}" is a non-decreasing we have "a_n>M" for all "n>N."

Finally: for all M>0 there exists "N\in\mathbb{N}" such that for all "n>N" "a_n>M". Therefore, "\{a_n\}_{n=1}^{+\infty}"diverges to +∞ by the definition.


2) Let "\{a_n\}_{n=1}^{+\infty}" is a non-inreasing sequence which is not bounded below, that is, "\inf\limits_{n}a_n = -\infty". This means that for all M<0 there exists "N\in\mathbb{N}" such that "a_N<M". As the sequence "\{a_n\}_{n=1}^{+\infty}" is a non-increasing we have "a_n<M" for all "n>N."

Finally: for all M<0 there exists "N\in\mathbb{N}" such that for all "n>N" "a_n<M". Therefore, "\{a_n\}_{n=1}^{+\infty}"diverges to -∞ by the definition.